A balloon was filled to 1.25L at 25.2°C. 0.235g of helium was used to fill it. What is the pressure?

R=0.8206L·atm/°C·mole
Answer #1

Applying, ideal gas equation, PV= nRT

Here ,n = no. of moles = 0.235/4.0 =0.05875 mole ( Molar mass of helium =4)

V= 1.25 L

T= 25.2°c

P=n/V *RT

= 0.05875mole/1.25 Litre * 0.8206 L.atm/°C.mole*25.2°c

=0.047*0.8206*25.2 atm

=0.9719 atm.

Thus, pressure = 0.9719 atm.

please upvote